# On estimation and order selection for multivariate extremes via clustering

### v1

@@ 1 Introduction
In this work, we further explore clustering-based estimation of multivariate extreme models with a discrete spectral measure. The contributions of this work are threefold. The main contribution involves the development of an order selection method that, on the theoretical side, consistently recovers the true order, and on the practical side, enjoys intuitive and simple implementation. Our method is based on a variant of the well-known Silhoutte method [25, 17]. In particular, we introduce an additional penalty term to the so-called simplified average silhouette width, which discourages small cluster sizes and small dissimilarity between cluster centers. As a consequence, we provide a method to consistently estimate the order of a max-linear factor model, for which a usual information-based method is not applicable due to the unavailability of likelihood (e.g., [5, 28]). Our second contribution concerns a large-deviation-type result on the discrete spectral measure estimation via the clustering methods such as the spherical $k$-means and $k$-pc. This constitutes an attempt to address the quality of convergence for clustering-based estimation in the context of multivariate extremes. As a third contribution, we discuss how the discrete spectral measure estimation can be translated into parameter estimates of the heavy-tailed max-linear and sum-linear factor models. Simulation and real-data studies illustrating order selection and factor model estimation are also provided.
The paper is organized as follows. Section 2 provides background and preliminary results on multivariate extremes, spherical clustering, and their connection. The penalized silhouette method for order selection is introduced in Section 3. Section 4 offers some large-deviation-type analysis of convergence of clustering-based spectral estimation. Section 5 relates clustering-based spectral estimation to the estimation of certain heavy-tailed factor models. Section 6 presents simulation and real-data demonstrations of order selection and factor model estimation. By default, all vectors are column vectors. The notation $\delta_{\mathbf{w}}$ stands for a delta measure with unit mass at the point $\mathbf{w}$ in an appropriate measurable space.
@@ 2.1 Multivariate extreme value theory
Suppose that $\mathbf{X}=(X_{1},\ldots,X_{d})$ is a $d$-dimensional random vector taking values in $[0,\infty)^{d}$ with continuous marginal distributions, where $d\geq 2$. Many discussions in this paper can be extended to the case of $\mathbb{R}^{d}$-valued $\mathbf{X}$, although for simplicity, we restrict ourselves to the nonnegative orthant $[0,\infty)^{d}$, which is also most commonly encountered in practice. As a conventional practice in the analysis of multivariate extremes, the modeling of marginals and extremal dependence is often separate. We assume that $\mathbf{X}$ has been marginally standardized to share a standard $\alpha$-Pareto-like tail asymptotically:
@@
where $\alpha>0$ is known, often chosen as $\alpha=1$ or $\alpha=2$ in literature. The so-called multivariate regular variation (MRV) assumption on $\mathbf{X}$ requires
@@
where $[\mathbf{0},\mathbf{x}]=[0,x_{1}]\times\ldots[0,x_{d}]$, $x_{i}>0$, $i\in\left\{1,\ldots,d\right\}$. Moreover, the measure $\Lambda$ satisfies the homogeneity property $\Lambda(c\,\cdot)=c^{-\alpha}\Lambda(\cdot)$, and therefore admits a polar decomposition into a product of radial and angular parts. We shall follow the formulation in [1, Section 8.2.5], which allows the use of different norms for the radial and angular components. Suppose that $\|\cdot\|_{(r)}$ and $\|\cdot\|_{(s)}$ denote two arbitrary norms on $\mathbb{R}^{d}$. Slightly abusing the notation, we still use $\Lambda$ denote the push-forward measure of $\Lambda$ under the one-to-one mapping $\mathbb{E}_{d}\mapsto(0,\infty)\times\mathbb{S}_{+}^{d-1}$, $\mathbf{x}\mapsto(r,\mathbf{w}=(w_{1},\ldots,w_{d})):=\left(\|\mathbf{x}\|_{(r)},\mathbf{x}/\|\mathbf{x}\|_{(s)}\right)$, where
@@
$\Lambda(dr,d\mathbf{w})=\Lambda_{\alpha}(dr,d\mathbf{w})=c_{(r)}\alpha r^{-\alpha-1}dr\times H(d\mathbf{w}),$ (5)
@@
In practice, commonly used norms are the $p$-norm $\|\mathbf{x}\|_{p}=\left(\sum_{j=1}^{d}|x_{j}|^{p}\right)^{1/p}$, $p\in(0,\infty)$, and the sup-norm $\|\mathbf{x}\|_{\infty}=\max\left(|x_{1}|,\ldots,|x_{d}|\right)$. In addition, the following weak convergence on $\mathbb{S}_{+}^{d-1}$ holds
@@ 2.2 Spherical clustering
The spherical clustering algorithms that have been considered so far are performed exclusively on the unit sphere $\mathbb{S}_{+}^{d-1}$ with respect to the $2$-norm (Euclidean norm), that is, take $\|\cdot\|_{(s)}$ in (4) as $\|\cdot\|_{2}$. We do not make this assumption for generality unless discussing specific examples. We equip $\mathbb{S}_{+}^{d-1}$ with the subspace topology inherited from $\mathbb{R}^{d}$. Next, we introduce a dissimilarity measure $D$ that follows the assumption below.
Assumption 1.
Suppose $D:\mathbb{S}_{+}^{d-1}\times\mathbb{S}_{+}^{d-1}\rightarrow[0,1]$ is continuous, and satisfies the following properties: for $\mathbf{w}_{i}\in\mathbb{S}_{+}^{d-1}$, $i\in\left\{1,2\right\}$, (i) $D(\mathbf{w}_{1},\mathbf{w}_{2})=0$ if and only if $\mathbf{w}_{1}=\mathbf{w}_{2}$; (ii) $D(\mathbf{w}_{1},\mathbf{w}_{2})=D(\mathbf{w}_{2},\mathbf{w}_{1})$.
@@
$B(\mathbf{w},r):=\{\mathbf{u}\in\mathbb{S}_{+}^{d-1}:\ D(\mathbf{w},\mathbf{u})<r\},$
@@
To simplify the mathematical description of clustering of sample data, it is convenient to use the notion of multiset. Recall that a multiset $W$ on $\mathbb{S}_{+}^{d-1}$ is a set that allows repetition of its elements, whose support, denoted as $\operatorname{supp}{W}$, is a subset of $\mathbb{S}_{+}^{d-1}$ in the usual sense that eliminates repetitions in $W$. For instance, with two distinct points $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$ on $\mathbb{S}_{+}^{d-1}$, one can have $W=\{\mathbf{w}_{1},\mathbf{w}_{1},\mathbf{w}_{2}\}$ with $\operatorname{supp}{W}=\{\mathbf{w}_{1},\mathbf{w}_{2}\}$. A multiset $W$ can be characterized by the multiplicity function $m_{W}:\mathbb{S}_{+}^{d-1}\mapsto\{0,1,2,\ldots\}$, where $m_{W}(\mathbf{w})$ equals the number of repetitions of element $\mathbf{w}\in\mathbb{S}_{+}^{d-1}$ ($m_{W}(\mathbf{w})=0$ if $\mathbf{w}\notin\operatorname{supp}{W}$). A subset of $\mathbb{S}_{+}^{d-1}$ in the usual sense can be understood as a multiset with the multiplicity taking value either 00 or $1$, with the empty set corresponding to a multiplicity function that is identically 00. When the notation $\mathbf{w}\in W$ is used for a multiset $W$, it means that $\mathbf{w}$ is an element in $\operatorname{supp}{W}$. For multisets $W_{1},W_{2}$ with multiplicity functions $m_{1}$ and $m_{2}$ respectively, their union $W_{1}\cup W_{2}$ is given by the multiset characterized by the multiplicity function $m_{1}\vee m_{2}$, and their intersection $W_{1}\cap W_{2}$ is given by the multiset characterized by $m_{1}\wedge m_{2}$. The relation $W_{1}\subset W_{2}$ is understood as $m_{1}\leq m_{2}$. Furthermore, if $\operatorname{supp}{W}$ is a finite set, a summation $\sum_{\mathbf{w}\in W}f(\mathbf{w})$ for a suitable function $f$ is understood as $\sum_{\mathbf{w}\in\operatorname{supp}{W}}f(\mathbf{w})m_{W}(\mathbf{w})$. For example, the cardinality of $W$ is defined as
@@
A $k$-clustering of a multiset $W$, $1\leq k\leq|W|$, with respect to the dissimilarity measure $D$ refers to a pair $(A_{k}^{*},\mathfrak{C}_{k})$. Here $A_{k}^{*}$ is as described above, and $\mathfrak{C}_{k}=\{C_{1},\ldots,C_{k}\}$ is a partition of $W$ into a collection of multisets $C_{i}$’s such that $D\left(\mathbf{w},A_{k}^{*}\right)=D\left(\mathbf{w},\mathbf{a}_{i}\right)$ for all $\mathbf{w}\in C_{i}$, $i\in\left\{1,\ldots,k\right\}$. We refer to $A_{k}^{*}$ as the set of centers and each $C_{i}$ as a cluster.
@@
With the choices $D=D_{\cos}$ and $D_{\rm{pc}}$ in (11) and (12), respectively, a $k$-clustering corresponds to the spherical $k$-means and $k$-pc clustering of [3] and [7], respectively. Solving a $k$-clustering problem can be computationally hard, and typically, the solution can only be approximated by a heuristic algorithm such as a Lloyd-type iterative algorithm as in [3] and [7]. In the theoretical analysis of this paper, we assume that a $k$-clustering can be found accurately. In addition, when $W$ is later given by a random subsample $W_{n}$ of the total sample $(\mathbf{X}_{i})_{i=1,\ldots,n}$, we assume that the elements in $A_{k}^{*}$ and the labels $\mathbf{1}\left\{\mathbf{X}_{i}\in C_{j}\right\}$, $i\in\{1,\ldots,n\}$, $j\in\{1,\ldots,k\}$, are measurable.
@@ 2.3 Spherical clustering for multivariate extremes
For any $S$ that is a $H$-continuity Borel subset of $\mathbb{S}_{+}^{d-1}$, we have $H_{n}(S)\rightarrow H(S)$ almost surely as $n\rightarrow\infty$.
@@
Now we consider applying the $k$-clustering in Definition 2.2 to the random subsample ${W}_{n}$. In particular, suppose that ${A}_{k,n}=\left({\mathbf{a}}_{1,n}^{k},\ldots,{\mathbf{a}}_{k,n}^{k}\right)$ is a multiset on $\mathbb{S}_{+}^{d-1}$ with cardinality $k$ that ${C}_{i,n}^{k},i\in\left\{1,\ldots,k\right\}$, are multisets on $\mathbb{S}_{+}^{d-1}$, such that $\left({A}_{k,n},\mathfrak{C}_{k,n}=\left\{{C}_{1,n}^{k},\ldots,{C}_{k,n}^{k}\right\}\right)$ form a $k$-clustering of $W_{n}$.
@@
Suppose $\mathbf{X}$ has a spectral measure of the following form:
@@
where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\ldots+p_{k}=1$. Let
@@
The convergence of $\mathbf{a}_{\pi_{n}(i),n}^{k}$ follows from [20, Theorem 3.1] (stated as convergence in Hausdorff distance between ${A}_{k,n}$ and $\{\mathbf{a}_{1},\ldots,\mathbf{a}_{k}\}$) and Proposition 2.4 above; see also the discussion in [20, Section 4]). It remains to show the convergence of $p_{\pi_{n}(i),n}^{k}$, $i\in\{1,\ldots,k\}$. Set
@@
Fix $\epsilon\in(0,r_{A}/3)$. By what has been proved and the continuity of $D^{\dagger}$, at almost every outcome $\omega$ of the sample space $\Omega$, for $n$ sufficiently large, we have $D^{\dagger}(\mathbf{a}_{\pi_{n}(i),n}^{k},\mathbf{a}_{i})<\epsilon$ , $i\in\{1,\ldots,k\}$. Fix for now such an $\omega$ and let $n$ be sufficiently large (possibly depending on $\omega$). Then by the triangular inequality (10), we have $B(\mathbf{a}_{i},\epsilon)\subset B\left(\mathbf{a}_{\pi_{n}(i),n}^{k},2\epsilon\right)\subset B(\mathbf{a}_{i},3\epsilon)$, $i\in\{1,\ldots,k\}$. Note that $B\left(\mathbf{a}_{\pi_{n}(i),n}^{k},2\epsilon\right)\cap W_{n}$ are disjoint for $i\in\{1,\ldots,k\}$. So in view of Definition 2.2, we have $B(\mathbf{a}_{i},\epsilon)\cap W_{n}\subset B\left(\mathbf{a}_{\pi_{n}(i),n}^{k},2\epsilon\right)\cap W_{n}\subset C_{\pi_{n}(i),n}^{k}\subset\left(B(\mathbf{a}_{i},\epsilon)\cup\cap_{j\neq i}B(\mathbf{a}_{j},\epsilon)^{c}\right)\cap W_{n}$, and hence
$H_{n}\left(B(\mathbf{a}_{i},\epsilon)\right)\leq p_{\pi_{n}(i),n}^{k}\leq H_{n}\left(B(\mathbf{a}_{i},\epsilon)\cup\cap_{j\neq i}B(\mathbf{a}_{j},\epsilon)^{c}\right),\quad i\in\{1,\ldots,k\}.$
The conclusion then follows from Proposition 2.4 since both sides above converges almost surely to $p_{i}$ as $n\rightarrow\infty$. ∎
@@
Comparing [20, Proposition 3.3] with Proposition 2.4 and Corollary 2.5 here, we have chosen to work directly under the marginal standardization assumption in (1) and not to treat the empirical marginal transformations as in [20, Eq. (3.5)] for simplicity. Nevertheless, the consistency result of the order selection below (Theorem 3.1) can be extended to the setup of [20] based on the results there.

### v2

@@ 1 Introduction
In this work, we further explore clustering-based estimation of multivariate extreme models with a discrete spectral measure. The contributions of this work are threefold. The main contribution involves the development of an order selection method that, on the theoretical side, consistently recovers the true order, and on the practical side, enjoys intuitive and simple implementation. Our method is based on a variant of the well-known silhouette method [29, 20]. In particular, we introduce an additional penalty term to the so-called simplified average silhouette width, which discourages small cluster sizes and small dissimilarity between cluster centers. As a consequence, we provide a method to consistently estimate the order of a max-linear factor model, for which a usual information-based method is not applicable due to the unavailability of likelihood (e.g., [7, 34]). Our second contribution concerns a large-deviation-type result on the discrete spectral measure estimation via the clustering methods such as the spherical $k$-means and $k$-pc. This constitutes an attempt to address the quality of convergence for clustering-based estimation in the context of multivariate extremes. As a third contribution, we discuss how the discrete spectral measure estimation can be translated into parameter estimates of the heavy-tailed max-linear and sum-linear factor models. Simulation and real-data studies illustrating order selection and factor model estimation are also provided.
The paper is organized as follows. Section 2 provides background and preliminary results on multivariate extremes, spherical clustering, and their connection. The penalized silhouette method for order selection is introduced in Section 3. Section 4 offers some large-deviation-type analysis of convergence of clustering-based spectral estimation. Section 5 relates clustering-based spectral estimation to the estimation of certain heavy-tailed factor models. Section 6 presents simulation and real-data demonstrations of order selection and factor model estimation. Section 7 provides a summary and a discussion on potential future directions. By default, all vectors are column vectors. The notation $\delta_{\mathbf{w}}$ stands for a delta measure with unit mass at the point $\mathbf{w}$ in an appropriate measurable space.
@@ 2.1 Multivariate extreme value theory
Suppose that $\mathbf{X}=(X_{1},\ldots,X_{d})$ is a $d$-dimensional random vector taking values in $[0,\infty)^{d}$, where $d\geq 2$. Many discussions in this paper can be extended to the case of $\mathbb{R}^{d}$-valued $\mathbf{X}$, although for simplicity, we restrict ourselves to the nonnegative orthant $[0,\infty)^{d}$, which is also most commonly encountered in practice. As a conventional practice in the analysis of multivariate extremes, the modeling of marginals and extremal dependence is often separate. We assume that $\mathbf{X}$ has been marginally standardized to share a standard $\alpha$-Pareto-like tail asymptotically:
@@
where $\alpha>0$ is known, often chosen as $\alpha=1$ or $\alpha=2$ in literature. One may generalize (1) to include slowly varying functions (see [2]), whereas we choose not to do so for simplicity. The so-called multivariate regular variation (MRV) assumption on $\mathbf{X}$ requires
@@
where $[\mathbf{0},\mathbf{x}]=[0,x_{1}]\times\ldots\times[0,x_{d}]$, $x_{i}>0$, $i\in\left\{1,\ldots,d\right\}$. Moreover, the measure $\Lambda$ satisfies the homogeneity property $\Lambda(c\,\cdot)=c^{-\alpha}\Lambda(\cdot)$, $c>0$, and therefore admits a polar decomposition into a product of radial and angular parts. We shall follow the formulation in [1, Section 8.2.5], which allows the use of different norms for the radial and angular components. Suppose that $\|\cdot\|_{(r)}$ and $\|\cdot\|_{(s)}$ denote two arbitrary norms on $\mathbb{R}^{d}$. Slightly abusing the notation, we still use $\Lambda$ denote the push-forward measure of $\Lambda$ under the one-to-one mapping $\mathbb{E}_{d}\mapsto(0,\infty)\times\mathbb{S}_{+}^{d-1}$, $\mathbf{x}\mapsto(r,\mathbf{w}=(w_{1},\ldots,w_{d})):=\left(\|\mathbf{x}\|_{(r)},\mathbf{x}/\|\mathbf{x}\|_{(s)}\right)$, where
@@
$\Lambda(dr,d\mathbf{w})=c_{(r)}\alpha r^{-\alpha-1}dr\times H(d\mathbf{w}),$ (5)
@@
In practice, commonly used norms are the $p$-norm $\|\mathbf{x}\|_{p}=\left(\sum_{j=1}^{d}|x_{j}|^{p}\right)^{1/p}$, $p\in(0,\infty)$, and the sup-norm $\|\mathbf{x}\|_{\infty}=\max\left(|x_{1}|,\ldots,|x_{d}|\right)$. In view of (2), the following weak convergence on $\mathbb{S}_{+}^{d-1}$ holds
@@ 2.2 Spherical clustering
The spherical clustering algorithms that have been considered so far are performed exclusively on the unit sphere $\mathbb{S}_{+}^{d-1}$ with respect to the $2$-norm (Euclidean norm), that is, take $\|\cdot\|_{(s)}$ in (4) as $\|\cdot\|_{2}$. We do not make this assumption for generality unless discussing specific examples. We equip $\mathbb{S}_{+}^{d-1}$ with the subspace topology inherited from $\mathbb{R}^{d}$. Next, we introduce a dissimilarity measure $D$ on $\mathbb{S}_{+}^{d-1}$ that is essential for clustering.
Definition 2.1.
A dissimilarity measure $D$ on $\mathbb{S}_{+}^{d-1}$ is a bivariate function $D:\mathbb{S}_{+}^{d-1}\times\mathbb{S}_{+}^{d-1}\rightarrow[0,1]$ that is continuous and satisfies the following properties: for $\mathbf{w}_{i}\in\mathbb{S}_{+}^{d-1}$, $i\in\left\{1,2\right\}$, (i) $D(\mathbf{w}_{1},\mathbf{w}_{2})=0$ if and only if $\mathbf{w}_{1}=\mathbf{w}_{2}$; (ii) $D(\mathbf{w}_{1},\mathbf{w}_{2})=D(\mathbf{w}_{2},\mathbf{w}_{1})$.
@@
$B_{D}(\mathbf{w},r):=\{\mathbf{u}\in\mathbb{S}_{+}^{d-1}:\ D(\mathbf{w},\mathbf{u})<r\},$
@@
Additionally, we have the following property: if $D(\mathbf{w}_{n},\mathbf{w})\to 0$ as $n\rightarrow\infty$, then $D^{\dagger}(\mathbf{w}_{n},\mathbf{w})\to 0$ as well.
@@
To simplify the mathematical description of clustering of sample data, it is convenient to use the notion of multiset (e.g., [24]). Recall that a multiset $W$ on $\mathbb{S}_{+}^{d-1}$ is a set that allows repetition of its elements, whose support, denoted as $\operatorname{supp}{W}$, is a subset of $\mathbb{S}_{+}^{d-1}$ in the usual sense that eliminates repetitions in $W$. For instance, with two distinct points $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$ on $\mathbb{S}_{+}^{d-1}$, one can have $W=\{\mathbf{w}_{1},\mathbf{w}_{1},\mathbf{w}_{2}\}$ with $\operatorname{supp}{W}=\{\mathbf{w}_{1},\mathbf{w}_{2}\}$. A multiset $W$ can be characterized by the multiplicity function $m_{W}:\mathbb{S}_{+}^{d-1}\mapsto\{0,1,2,\ldots\}$, where $m_{W}(\mathbf{w})$ equals the number of repetitions of element $\mathbf{w}\in\mathbb{S}_{+}^{d-1}$ ($m_{W}(\mathbf{w})=0$ if $\mathbf{w}\notin\operatorname{supp}{W}$). A subset of $\mathbb{S}_{+}^{d-1}$ in the usual sense can be understood as a multiset with the multiplicity taking value either 00 or $1$, with the empty set corresponding to a multiplicity function that is identically 00. When the notation $\mathbf{w}\in W$ is used for a multiset $W$, it means that $\mathbf{w}$ is an element in $\operatorname{supp}{W}$. For multisets $W_{1},W_{2}$ with multiplicity functions $m_{1}$ and $m_{2}$ respectively, their union $W_{1}\cup W_{2}$ is given by the multiset characterized by the multiplicity function $m_{1}\vee m_{2}$, and their intersection $W_{1}\cap W_{2}$ is given by the multiset characterized by $m_{1}\wedge m_{2}$. The relation $W_{1}\subset W_{2}$ is understood as $m_{1}\leq m_{2}$. Furthermore, if $\operatorname{supp}{W}$ is a finite set, a summation $\sum_{\mathbf{w}\in W}f(\mathbf{w})$ for a suitable function $f$ is understood as $\sum_{\mathbf{w}\in\operatorname{supp}{W}}f(\mathbf{w})m_{W}(\mathbf{w})$. For example, the cardinality of $W$ is defined as
@@
A $k$-clustering of a multiset $W$, $1\leq k\leq|W|$, with respect to the dissimilarity measure $D$ refers to a pair $(A_{k}^{*},\mathfrak{C}_{k})$. Here $A_{k}^{*}=\left\{\mathbf{a}_{1}^{*},\ldots,\mathbf{a}_{k}^{*}\right\}$ is as described above, and $\mathfrak{C}_{k}=\{C_{1},\ldots,C_{k}\}$ is a partition of $W$ into a collection of multisets $C_{i}$’s such that $D\left(\mathbf{w},A_{k}^{*}\right)=D\left(\mathbf{w},\mathbf{a}_{i}^{*}\right)$ for all $\mathbf{w}\in C_{i}$, $i\in\left\{1,\ldots,k\right\}$. We refer to $A_{k}^{*}$ as the set of centers and each $C_{i}$ as a cluster.
@@
With the choices $D=D_{\cos}$ and $D_{\rm{pc}}$ in (11) and (12), respectively, a $k$-clustering corresponds to the spherical $k$-means and $k$-pc clustering of [6] and [10], respectively. Solving a $k$-clustering problem can be computationally hard, and typically, the solution can only be approximated by a heuristic algorithm such as a Lloyd-type iterative algorithm as in [6] and [10]. In the theoretical analysis of this paper, we assume that a $k$-clustering can be found accurately. In addition, when $W$ is a random multiset, for a $k$-clustering of $W$, we assume that the elements in $A_{k}^{*}$ and the labels $\mathbf{1}\left\{\mathbf{w}\in C_{j}\right\}$, $\mathbf{w}\in W$, $j\in\{1,\ldots,k\}$, are measurable.
@@ 2.3 Spherical clustering for multivariate extremes
Suppose $\mathbf{X}$ satisfies conditions (1) and (2) with a spectral measure $H$ on $\mathbb{S}_{+}^{d-1}$ as defined in (5). Let $W_{n}$ denote the extremal subsample as in (14) and $H_{n}$ denote the empirical spectral measure as in (17). Then for any $S$ that is a $H$-continuity Borel subset of $\mathbb{S}_{+}^{d-1}$, we have $H_{n}(S)\rightarrow H(S)$ almost surely as $n\rightarrow\infty$.
@@
Now we consider applying the $k$-clustering in Definition 2.3 to the random subsample ${W}_{n}$. In view of Proposition 2.5, the distribution of $W_{n}$ on $\mathbb{S}_{+}^{d-1}$ serves as a good approximation to $H$. When $H$ is a discrete measure with finitely many atoms, clustering can be applied to $W_{n}$ in such a way that the cluster centers closely approximate the locations of the atoms, and the cluster proportions provide accurate estimates of the associated probabilities. Furthermore, these estimators are shown to be consistent, as demonstrated in the following Corollary.
@@
Suppose $\mathbf{X}$ is as in Proposition 2.5, and furthermore, has a spectral measure of the following form:
@@
where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\ldots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14), and $\left({A}_{k,n}=\left({\mathbf{a}}_{1,n}^{k},\ldots,{\mathbf{a}}_{k,n}^{k}\right),\mathfrak{C}_{k,n}=\left\{{C}_{1,n}^{k},\ldots,{C}_{k,n}^{k}\right\}\right)$ form a $k$-clustering of $W_{n}$ as defined in Definition 2.3 with respect to a dissimilarity measure $D$ defined in Definition 2.1. Let
@@
The convergence of $\mathbf{a}_{\pi_{n}(i),n}^{k}$ follows from [23, Theorem 3.1] (stated as convergence in Hausdorff distance between ${A}_{k,n}$ and $\{\mathbf{a}_{1},\ldots,\mathbf{a}_{k}\}$) and Proposition 2.5 above; see also the discussion in [23, Section 4]). It remains to show the convergence of $p_{\pi_{n}(i),n}^{k}$, $i\in\{1,\ldots,k\}$. Set
@@
Fix $\epsilon\in(0,r_{A}/3)$. By what has been proved and the continuity of $D^{\dagger}$, at almost every outcome $\omega$ of the sample space $\Omega$, for $n$ sufficiently large, we have the dual dissimilarity $D^{\dagger}(\mathbf{a}_{\pi_{n}(i),n}^{k},\mathbf{a}_{i})<\epsilon$ , $i\in\{1,\ldots,k\}$. Fix for now such an $\omega$ and let $n$ be sufficiently large (possibly depending on $\omega$). Then by the triangular inequality (10), we have $B_{D}(\mathbf{a}_{i},\epsilon)\subset B_{D}\left(\mathbf{a}_{\pi_{n}(i),n}^{k},2\epsilon\right)\subset B_{D}(\mathbf{a}_{i},3\epsilon)$, $i\in\{1,\ldots,k\}$. Note that $B_{D}\left(\mathbf{a}_{\pi_{n}(i),n}^{k},2\epsilon\right)\cap W_{n}$ are disjoint for $i\in\{1,\ldots,k\}$ due to the choice of $\epsilon$. So in view of Definition 2.3, we have $B_{D}(\mathbf{a}_{i},\epsilon)\cap W_{n}\subset B_{D}\left(\mathbf{a}_{\pi_{n}(i),n}^{k},2\epsilon\right)\cap W_{n}\subset C_{\pi_{n}(i),n}^{k}\subset\left(B_{D}(\mathbf{a}_{i},\epsilon)\cup\cap_{j\neq i}B_{D}(\mathbf{a}_{j},\epsilon)^{c}\right)\cap W_{n}$, and hence
$H_{n}\left(B_{D}(\mathbf{a}_{i},\epsilon)\right)\leq p_{\pi_{n}(i),n}^{k}\leq H_{n}\left(B_{D}(\mathbf{a}_{i},\epsilon)\cup\cap_{j\neq i}B_{D}(\mathbf{a}_{j},\epsilon)^{c}\right),\quad i\in\{1,\ldots,k\}.$
Recall the inequality above is derived for almost every $\omega$ with $n$ sufficiently large. The conclusion then follows from Proposition 2.5 since both sides above converges almost surely to $p_{i}$ as $n\rightarrow\infty$. ∎
@@
Comparing [23, Proposition 3.3] with Proposition 2.5 and Corollary 2.6 here, we have chosen to work directly under the marginal standardization assumption in (1) and not to treat the empirical marginal transformations as in [23, Eq. (3.5)] for simplicity. Nevertheless, the consistency result of the order selection below (Theorem 3.1) can be extended to the setup of [23] based on the results there.